### v1

@@ V-B Motion capture system
To provide a ground-truth comparison for the sensors, we used a twelve-camera OptiTrack Flex 13 motion capture arena. The Crazyflie drone was fitted with four retro-reflective motion capture markers and its position and orientation were recorded at 120 Hz. Velocity was estimated by taking the numerical derivative of the position data in post-processing.
@@ V-C Experiment
Figure 5: State estimates from TinySense after Kalman filtering, Crazyflie estimates, and mocap estimates are compared. We conducted three flight experiments (a), (b), and (c), with the Kalman filter estimates starting at time zero. Note that we ignore the pressure sensor measurements in the first 1.8s because of the pressure change near the surface.
TABLE II: Estimation of Optic Flow Algorithm Power Use
Update Frequency (Hz) Occurences/ Update Step Name Single-cycle Operations Int Division Operations Floating-point Division Operations Total Cycles/Update Total Cycles/Second Power Use
100 1 Skip frame 8493 0 8 249781 24.9781 Mhz $\sim$5.4552mW
@@ V-D Computational Load
We estimated the power consumption of the optic flow algorithm from the number of cycles in a single run. We focus on only estimating three types of operations conducted on the ARM Cortex-M4 based microcontroller: single-cycle operations (addition, subtraction, multiplication, and memory read/write); integer divisions, which require 12 cycles at most; and floating-point divisions, which require 14 cycles. These results are summarized in Table II. Given that in our current system, all sensors update at a rate of 100 Hz, we gather that the total number of cycles per second is $\sim$24.98 Mhz. Operating on the nRF52840 chip with an efficiency of 52 $\mu$A/Mhz and a battery-powered voltage of 4.2V, we estimate the power use of the optic flow algorithm to be $\sim$5.45mW.
V-E Results and Discussions
Fig. 5 compares our sensor suite’s state estimate to that of the Crazyflie’s built-in estimator, and states derived from the mocap system, considered to be ground truth. For the state estimator, time $t=0$ corresponds to when the gyroscope reading becomes significantly different from zero, indicating liftoff. It can be observed in Fig. 4 that a pressure increase occurs when the propellers first turn on, which can be observed as a drop in measured altitude. To mitigate unwanted effects of this in the Kalman Filter, we ignored (set to zero) altitude measurements $z_{m}$ during the first 1.8s after liftoff. Nevertheless, once it is in the air, the altitude measurement matches quite closely with the laser rangefinder on the Crazyflie, though with slightly more noise.
Our system’s estimate compares well with those of the Crazyflie’s, despite receiving gyroscope readings at a much lower rate (100 Hz) compared to 1000 Hz on the Crazyflie[35]. The RMSE is approximately equal for pitch angle (Table III). We observe that there are occasionally large errors in translational velocity $v_{x}$. These appear to occur after sudden changes in attitude that occur around $t=5$ s and $t=6$ s, during which the angular velocity momentarily exceeds 3 rad/s (Fig. 4). We note that at these moments, our TinySense optic flow sensor’s readings are consistently lower than those of the Crazyflie. Lucas-Kanade cannot detect motion accurately if the displacement between frames is greater than a half pixel between image frames. The maximum allowable angular velocity is given by $\Omega=0.5\alpha F=(0.5)(0.027)(100)=1.4$ rad/s, where $\alpha$ is the angular increment between neighboring pixels (Section II-A) and $F$ is the frame rate in Hz. So the expected optic flow is faster than can be detected. This could be mitigated by a faster frame rate or a larger angular displacement between pixels.
@@ VI Conclusion
In the present study, we introduce a new sensor suite with dramatically reduced mass, and associated optic flow and state estimation software that is compatible with the computational constraints of an onboard microcontroller small enough to fly aboard flying insect robots (FIRs) weighing less than a gram. Our proposed avionics suite diverges in two important ways compared to that of the 28 g Crazyflie helicopter, the lightest vehicle yet to perform controlled hovering, and previous work [15]. First, it replaces a power-hungry laser rangefinder with a much more efficient (and slightly less ideal) and lighter pressure sensor. Second, we replaced an off-the shelf optic flow sensor, which is available only in relatively heavy packages, with a custom camera with a global shutter and a custom-written Lucas-Kanade-based optic flow algorithm running onboard a 10 mg microcontroller small enough to fly onboard an FIR (estimated conpute power usage is about 5 mW). Our system uses a Kalman Filter to estimate pitch angle, translational velocity, and altitude. Remarkably, the cumulative weight of this sensor suite is only 78.4 mg, rather less than the 187 mg of previous work [15]. Incorporating a microcontroller to this configuration, the overall system remains well within the anticipated 252 mg payload limit of the 143 mg robotic platform reported in [18]. Our system demonstrates average RMSE values of 0.186 m/s for velocity, 1.573 degrees for pitch angle, and 0.139 meters for altitude when compared to the motion capture system. In comparison, the Crazyflie platform achieves RMSE values of 0.075 m/s, 1.619 degrees, and 0.021 meters for the same states during similar flight maneuvers. This shows that our system performs closely to the state-of-the-art Crazyflie in these scenarios. In future work, we aim to integrate TinySense with insect-scale robots to perform more advanced tasks, such as confined-space navigation [36].

### v2

@@ V-B Motion capture system
To provide a ground-truth comparison for the sensors, we used a twelve-camera OptiTrack Flex 13 motion capture arena, with spatial accuracy $\pm$ 0.20 mm. The Crazyflie drone was fitted with four retro-reflective motion capture markers and its position and orientation were recorded at 120 Hz. Velocity was estimated by taking the numerical derivative of the position data in post-processing.
@@ V-C Experiment
Figure 5: Comparison between state estimates from TinySense’s Kalman Filter, Crazyflie, and mocap. We conducted three flight experiments (a), (b), and (c), with the Kalman filter estimates starting at time zero. Pressure sensor measurements were ignored during the first 1.8 s because pressure changes due to the ground effect at takeoff confound readings.
TABLE II: Estimate of optic flow algorithm power use
update frequency (Hz) occurences/ update task single-cycle operations int division operations floating-point division operations total cycles/update total cycles/s (Mhz) estimated power (mW)
100 1 skip frame 8493 0 8 249781 25.0 5.5
@@ V-D Computational load
Table II provides an estimate of the power consumption of the optic flow algorithm, by calculating the number of clock cycles needed for each operation. Our estimate includes the matrix operations described in Section III-A, as well as additional housekeeping operations, and is based on ARM Cortex-M4 instruction timings for addition, subtraction, multiplication, and memory read/write operations; integer divisions require 12 cycles at most and floating-point divisions require 14 cycles. Given that in our current system all sensors update at a rate of 100 Hz, we estimated that the total number of needed cycles per second is $\sim$25.0 Mhz. Operating the nRF52840 with an efficiency of 52 $\mu$A/MHz and a battery-powered voltage of 4.2 V, we estimate the power usage of the optic flow algorithm to be 5.5 mW.
V-E Results and discussion
Fig. 5 compares our sensor suite’s state estimate to that of the Crazyflie’s built-in estimator. States derived from the mocap system are considered to be ground truth. For the state estimator, time $t=0$ corresponds to when the gyroscope reading becomes significantly different from zero, indicating liftoff. Fig. 4 shows that a pressure increase occurs when the propellers first turn on, which can be observed as a drop in measured altitude. To mitigate unwanted effects of this in the Kalman Filter, we ignored (set to zero) altitude measurements $z_{m}$ during the first 1.8 s after liftoff. Nevertheless, once it is in the air, the altitude measurement matches quite closely with the laser rangefinder on the Crazyflie, though with slightly more noise.
Our system’s estimate compares well with the Crazyflie’s, despite receiving gyroscope readings at only 100 Hz, which is ten times slower [37]. The key metric of pitch angle RMSE is approximately equal (Table III). Large errors in translational velocity $v_{x}$ appear around $t=5$ and $t=6$. These occur after sudden changes in attitude, during which angular velocity exceeds 3 rad/s (Fig. 4). At these moments, we observed that TinySense optic flow readings are consistently lower than those of the Crazyflie. Lucas-Kanade cannot detect motion accurately if the displacement between frames is greater than a half pixel between image frames [38]. This limits optic flow measurement to $\Omega_{m}<0.5\alpha F=1.4$ rad/s, where $\alpha$=0.027 is the angular increment between pixels and $F$=100 is the frame rate (Hz). We believe that the lateral velocity estimate diverges during these fast maneuvers because it depends on erroneous optic flow readings.
@@ VI Conclusion
In the present study, we introduce a new sensor suite with dramatically reduced mass, and associated optic flow and state estimation software that is compatible with the computational constraints of an onboard microcontroller small enough to fly aboard flying insect robots (FIRs) weighing less than a gram. Our proposed avionics suite diverges in two important ways from the 28 g Crazyflie helicopter, the lightest vehicle yet to perform controlled hovering, and previous work [15]. First, we replaced a power-hungry laser rangefinder with a much more efficient (and slightly less ideal) and lighter pressure sensor. Second, we replaced an off-the-shelf optic flow sensor, which is available only in relatively heavy packages, with a custom camera with a global shutter and a custom-written Lucas-Kanade-based optic flow algorithm running onboard a 10 mg microcontroller small enough to fly onboard an FIR. We estimate the compute power usage is about 5 mW. Our system uses a linear Kalman Filter to estimate pitch angle, translational velocity, and altitude. This can be readily adapted to a nonlinear extended Kalman Filter for more generality [15].
The cumulative weight of this sensor suite is only 78.4 mg, rather less than the 187 mg of previous work [15]. With the addition of a 30 mg microcontroller and crystal, the overall system remains well within the 252 mg payload limit of the 143 mg robotic platform reported in [20]. Our system demonstrates average RMSE values of 0.186 m/s for velocity, 1.573 deg for pitch angle, and 0.136 m for altitude when compared to the motion capture system. In comparison, the Crazyflie platform achieves RMSE values of 0.075 m/s, 1.619 deg, and 0.021 m for the same states during similar flight maneuvers. This shows that our system performs closely to the state-of-the-art Crazyflie in these scenarios. The pressure sensor showed negligible drift over a 60-second period, providing a usable altitude estimate as long as its output during takeoff is ignored. Future improvements include mitigating the lateral velocity errors that occur during rapid rotations that exceed the limit of Lucas-Kanade estimation (Section V-E) by increasing the frame rate or image blurring. Additionally, our patch-based optic flow estimation (Section III-A) may allow outliers to be rejected based on their eigenvalues to eliminate low-contrast patches, reducing noise [38].
Future work will integrate TinySense along with a microcontroller and power hardware onto insect-scale robots such as the Robofly [39] to perform visual navigation [40]. The Robofly performs faster accelerations than the Crazyflie and is a different vibration environment, but is otherwise similar. Previous work shows vibrations do not confound gyroscope readings [41, 16]. Algorithmic changes in optic flow may be needed to accommodate vibration, but our choice of a global shutter camera will mitigate tearing artifacts.